OpenStudy (anonymous):

i need help solving a system from the points (0,4) and (1,3)

OpenStudy (icecreamyummy):

do you know how to find the slope

OpenStudy (icecreamyummy):

use www.mathways.com that's helps alot

OpenStudy (anonymous):

what are your x and y's

OpenStudy (icecreamyummy):

Xs are always first ys are second not mater what it is allways like that (X,Y)

OpenStudy (anonymous):

then count up how many spaces and over how many spaces to the next point. for the points (-3,4) and (2,-1) the slope would be 1/1

OpenStudy (icecreamyummy):

very good

OpenStudy (icecreamyummy):

what is the y intercept it is lie (0,___) the blank is a y intercept

OpenStudy (texaschic101):

are you trying to find the slope given points (0,4) and (1,3) ??

OpenStudy (texaschic101):

because it is not 1/1

OpenStudy (anonymous):

oops. looks Like I need help too.

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (0,4)...x1 = 0 and y1 = 4 (1,3)...x2 = 1 and y2 = 3 now sub slope(m) = (3 - 4) / (1 - 0) slope(m) = -1/1

OpenStudy (anonymous):

i was so close

OpenStudy (texaschic101):

now what exactly do you need to do for this question ? Do you need to come up with the line that these points are on ?

OpenStudy (texaschic101):

or do you just need to find x and y intercepts ?

OpenStudy (anonymous):

no, i need tofind the y intercept and then the x, and lastly i need to use subsitution to find where the lines intersect.

OpenStudy (texaschic101):

we need the equation use y = mx + b we already know slope(m) = -1...so we need to find y intercept (b) use either of your points (1,3)..x = 1 and y = 3 now sub y = mx + b 3 = -1(1) + b 3 = -1 + b 3 + 1 = b 4 = b so your equation of your line is : y = -1x + 4 are you still with me ?

OpenStudy (texaschic101):

are you needing to find the line for (-3,4) and (2,-1) ??